A solution is made by combining 200.0 mL of 0.23 mol/L HCl, 600.0 mL of 0.16mol/L KOH, and 200.0 mL of water. What is the pH of this solution?
1
Expert's answer
2020-10-02T14:12:26-0400

200ml of 0.23mol/L HCL =(volume*concentration)=(200*0.23)=46milimole HCL

600ml of 0.16mol/L KOH =(600*0.16)=96 milimole KOH,

The reaction between HCL and KOH is

HCL+KOH=KCL+H2O

46 milimole HCL react with 46 milimole KOH ,excess KOH in solution is (96-46)=50 milimole=50*10^-3 mole

Total volume of the solution is (200+600+200)ml =1000ml= 1L

Molar concentration of KOH is 50*10^-3(M)

pOH= -log(OH-)= -log(5*10^-3)=2.30

We know pH+pOH= 14

Or,pH = 14-pOH

pH= 14-2.30=11.7

Hence the pH of the resulting solution is 11.7